OpenStudy (anonymous):

what is not an example of exponential function. a.population growth b.exponential decay c.compund interest d.simple interest

OpenStudy (anonymous):

Simple interest is given by: \[SI = \frac{P \times R \times T}{100}\] Does that look exponential to you??

OpenStudy (anonymous):

no it doesnt

OpenStudy (anonymous):

Are you seeing any power in this ?? There is no power so it is not exponential.. Yes you are right...

OpenStudy (anonymous):

oh oay well thank you :)

OpenStudy (anonymous):

Welcome dear..